# NAVSIM: Data-Driven Non-Reactive Autonomous Vehicle Simulation and Benchmarking

### v1

@@ Abstract
Figure 1: NAVSIM. Traditional metrics such as the average displacement error (ADE) do not account for the interactive and multi-modal nature of driving. Our benchmark uses simulation-based metrics to provide more meaningful evaluations of trajectory outputs from sensor-based driving policies.
@@ 1 Introduction
Unfortunately, it is extremely challenging to evaluate driving performance, and the most widely-used benchmarks today fall short in several respects: (1) the datasets used, such as nuScenes [5], were created for perception tasks such as object detection. As such, they focus on visual diversity and label quality instead of the relevance of the data for research on planning. Often, most frames have a trivial solution of extrapolating the historical driving behavior, leading to “blind” driving policies that observe only the vehicle’s past trajectory obtaining state-of-the-art performance  [56, 32, 16]. (2) Due to the fact that driving is an inherently multifaceted task where the algorithm must balance several desired properties such as safety, comfort, and progress, the evaluation must also involve multiple complementary metrics. However, as shown in Fig. 1, existing metrics such as the average displacement error (ADE) between a predicted and recorded human trajectory often misrepresent the relative accuracy of trajectories. (3) Since driving involves interactions among multiple agents, evaluation must ideally be interactive, e.g., in simulation. Unfortunately, existing simulators with synthetic sensor data exhibit a significant domain gap to real-world driving. (4) Besides, the lack of a standardized evaluation setup has led to subtle inconsistencies between metrics in existing work, leading to unfair comparisons and inaccurate conclusions [50, 32]. Collectively, these problems hinder progress in the development of AVs, emphasizing the need for more principled benchmarks.
@@ 2 Related Work
Open-Loop Benchmarking with Displacement Errors. Open-loop evaluation protocols commonly measure displacement errors between trajectories of a recorded expert (i.e., of a human driver) and a motion planner. However, several issues concerning evaluation with displacement errors have surfaced recently, particularly on the nuScenes dataset [5]. Given that nuScenes does not provide standardized planning metrics, prior work relied on independent implementations, which led to inconsistencies when reporting or comparing results [50, 32]. Next, most planning models in nuScenes receive the human trajectory endpoint as a discrete direction command [20, 21, 25, 32, 50], thereby leaking ground-truth information into inputs. Moreover, about 75% of the scenarios in nuScenes involve straight driving [32], leading to simple solutions when extrapolating the ego-motion. For instance, AD-MLP demonstrates that an MLP on the kinematic ego status (ignoring perception completely) can achieve state-of-the-art displacement errors [56]. Such blind agents are undeniably dangerous, which highlights a broader concern: displacement metrics are not correlated to closed-loop driving [14, 17, 3, 16]. In this work, we address prevalent issues of nuScenes and propose a standardized driving benchmark with challenging scenarios and an official evaluation server. We derive a navigation goal from the lane graph instead of the human trajectory to prevent label leakage, and propose principled simulation-based metrics as an alternative to displacement errors.
@@ 3 NAVSIM: Non-Reactive Autonomous Vehicle Simulation
Task description. Driving agents in NAVSIM must plan a trajectory, defined as a sequence of future poses, over a horizon of four seconds. Their input contains streams of past frames from onboard sensors, such as cameras, LiDAR, as well as the vehicle’s current speed, acceleration, and navigation goal, jointly termed the ego status. For compatibility with prior work [20, 21, 25, 50], we provide the navigation goal as a one-hot vector with three categories: left, straight, or right.
Non-Reactive Simulation. Traditional closed-loop benchmarks normally infer planners at high frequencies (e.g., 10Hz) [17, 26]. However, this requires efficient simulation of all input modalities of the driving agent, including high-dimensional sensor streams in the case of sensor-based approaches. To sidestep this, the core idea of NAVSIM is to evaluate driving agents using a non-reactive simulation. This means driving agents are only employed in the initial frame of each scene. Afterwards, the planned trajectory is kept fixed for the entire trajectory duration. Over this short horizon, no environmental feedback is provided to the driving agent, and the NAVSIM evaluation is purely based on the initial real-world sensor sample. This makes the agent’s task more challenging, as it requires a safe plan up to the entire four-second horizon. Therefore, it is not possible to scale this kind of simulation to arbitrarily long time horizons. Despite this limitation, non-reactive simulation offers a key advantage: unlike traditional open-loop benchmarks, which mainly compare the planned trajectory to the human driver’s trajectory in a similar setting, it enables the use of simulation outcomes to compute metrics reflecting safety, comfort, and progress. An LQR controller [28] is applied at each simulation iteration to calculate steering and acceleration values, and a kinematic bicycle model [37] propagates the ego vehicle. We execute this pipeline at $10$Hz over the $4$s trajectory horizon. In Sec. 4.1, we show that despite our simplifying assumption, our evaluation results in a much better alignment with closed-loop metrics than traditional open-loop metrics achieve.
@@ 3.1 Generating Standardized and Challenging Train and Test Splits
Dataset. The NAVSIM framework is agnostic to the choice of driving dataset. We choose OpenScene [15], a redistribution of nuPlan [26], the largest annotated public driving dataset. OpenScene includes 120 hours of driving at a reduced frequency of $2$Hz typically considered by end-to-end planning algorithms, resulting in a 10×10$\times$10 × reduction of data storage requirements compared to nuPlan from over 20 TB to 2 TB. Our agent input, based on OpenScene, comprises eight cameras, each with a resolution of $1920\times 1080$ pixels, and a merged LiDAR point cloud from five sensors. The input includes the current time-step and optionally 3 past frames, totaling $1.5$s at $2$Hz. In principle, any driving dataset that provides annotated HD maps, object bounding boxes, and sensor data can be converted into this format and thus be used with NAVSIM.
@@ 4 Experiments
Figure 3: Closed-Loop Alignment. (a) For each planner, we show open-loop metrics (OLS, PDMS) together with the corresponding closed-loop score (CLS). The trendlines depicting correlations are fit linearly to all (learned and rule-based) planners. Moreover, we analyze different (b) CLS durations $d$, (c) planning frequencies $f$, (d) PDMS horizons $h$, and (e) background agent behaviors.
@@ 4.1 Alignment Between Open-Loop and Closed-Loop Evaluation
Due to the heavy computational requirements of closed-loop simulation, we evaluate on the navmini split. This is a new split we create for rapid testing, with 396 scenarios in total that are independent of both navtrain and navtest but filtered using the same strategy (Section 3.1) and hence similarly distributed. We note that nuPlan offers two kinds of background agents: reactive agents along lane centers based on the Intelligent Driver Model (IDM) [48], and non-reactive agents replayed from the dataset, which we employ unless otherwise stated. We also default to a closed-loop simulation duration of $d=15\textrm{s}$, and a planning frequency of $f=10\textrm{Hz}$, which are standard for nuPlan.
@@
Results. The alignment between metrics is presented in Fig. 3 (a-e). Compared to OLS, we consistently observe better closed-loop correlation for PDMS, in terms of Spearman’s (rank) and Pearson’s (linear) correlation coefficients. As shown in (a), PDMS can capture the closed-loop properties of both learned and rule-based planners, whereas distance-based open-loop metrics show a clear misalignment. Decreasing the CLS duration in (b) from $d=15$s to $d=4$s further raises the correlation of PDMS and OLS, as the simulation horizon more closely matches the open-loop counterparts. Interestingly, we observe a higher correlation of open-loop metrics in (c) when reducing the planning frequency to $2$Hz. We expect a lower planning frequency to mitigate cumulative errors and enhance the controller’s stability in simulation, leading to more precise trajectory execution. Moreover, we observe an increase in correlation for longer PDMS horizons in (d), ranging from $h=2$s to $h=8$s. While predicting the future motion over 8s is challenging in uncertain scenarios, our results indicate the value of long horizons when evaluating motion planners. Lastly, replacing the non-reactive background agents with reactive IDM vehicles in (e) has little effect on the correlation, possibly due to the similar difficulty of both tasks [16].
@@ 4.2 Analysis of the State of the Art in End-to-End Autonomous Driving
Methods. As a lower bound, we consider the (1) Constant Velocity baseline detailed in Section 3.1. We include an (2) Ego Status MLP as a second "blind" agent, which leverages an MLP for trajectory prediction given only the ego velocity, acceleration and navigation goal. As an established architecture on CARLA, we evaluate our reimplementation of (3) TransFuser [12], which uses three cropped and downscaled forward-facing cameras, concatenated into a $1024\times 256$ image, and a rasterized BEV LiDAR input for predicting waypoints. It performs 3D object detection and BEV semantic segmentation as auxiliary tasks. We then consider (4) Latent TransFuser (LTF) [12], which shares the same architecture as TransFuser but replaces the LiDAR input with a learned embedding, hence requiring only camera inputs. Moreover, we provide two state-of-the-art end-to-end architectures for open-loop trajectory prediction on nuScenes. (5) UniAD [21] incorporates a wide range of tasks, such as mapping, tracking, motion, and occupancy prediction in a semi-sequential architecture, which processes feature representations through several transformer decoders culminating in a trajectory planning module. (6) PARA-Drive [50] uses the same auxiliary tasks, but parallelizes the network architecture, and the auxiliary task heads are trained in parallel with a shared encoder. Both UniAD and PARA-Drive use a BEVFormer backbone [31], which encodes the eight surround-view $1920\times 1080$ camera images over four temporal frames into a BEV feature representation. Implementation details for all methods are provided in the supplementary material.
@@ 5 Discussion
We present NAVSIM, a framework for non-reactive AV simulation. We address several shortcomings of existing driving benchmarks and propose standardized but configurable simulation-based metrics for benchmarking driving policies. We improve accessibility for conducting AV research by providing downloadable challenging scenario splits and simple data curation methods. We demonstrate that our evaluation protocol is better aligned to closed-loop driving, benchmark an established set of end-to-end planning baselines, and present the results of our inaugural competition. We hope that NAVSIM can serve as an accessible toolkit for AV researchers that bridges the gap between simulated and real-world driving. We aim to support more datasets in the future, and advocate for more open dataset releases by the community for accelerating progress in autonomous driving.
Limitations. While we show improvements over displacement error-based benchmarking, several aspects of driving remain unaddressed by evaluation in NAVSIM. A high PDMS does not always imply a high CLS, since our framework does not consider reactiveness or compounding accumulation of errors in closed-loop simulation. Moreover, closed-loop metrics also face problems, i.e., PDMS inherits several weaknesses of nuPlan’s CLS. Both scores do not regard certain traffic rules (e.g., stop-sign or traffic light compliance) or concepts such as transit and fuel efficiency. Moreover, as in CLS, rear-end collisions into the ego vehicle are currently not classified as "at-fault", resulting in little importance given to the scene behind the vehicle in NAVSIM. Further, certain limitations of the nuPlan dataset persist in NAVSIM, such as minor errors in camera parameters or noise in poses and 3D annotations. Our analysis might favor methods that are robust to such inconsistencies. In the future, we aim to improve the subscore definitions (e.g. the at-fault collision logic) and add more subscores during aggregation. Additionally, we would like to explore pose or route augmentations to account for control drifts in closed-loop driving and improve our evaluation’s diversity. Given these limitations, we strongly encourage the use of graphics-based closed-loop simulators, such as CARLA [17], as complementary benchmarks to NAVSIM when developing planning algorithms.

### v2

@@ Abstract
Figure 1: NAVSIM. Traditional metrics such as the average displacement error (ADE) overlook the multi-modality of driving. They penalize trajectories that deviate from a recorded human driving log, even if such a trajectory is safe. Our benchmark evaluates trajectory outputs of sensor-based driving policies with simulation-based metrics, considering collisions and map compliance.
@@ 1 Introduction
Unfortunately, it is extremely challenging to evaluate driving performance, and the most widely-used benchmarks today fall short in several respects: (1) the datasets used, such as nuScenes [5], were created for perception tasks such as object detection. As such, they focus on visual diversity and label quality instead of the relevance of the data for research on planning. Often, most frames have a trivial solution of extrapolating the historical driving behavior, leading to “blind” driving policies that observe only the vehicle’s past trajectory obtaining state-of-the-art performance  [56, 32, 16]. (2) Due to the fact that driving is an inherently multifaceted task where the algorithm must coordinate several desired properties such as safety, comfort, and progress, the evaluation metric must also balance potentially conflicting objectives. However, as shown in Fig. 1, existing metrics such as the average displacement error (ADE) between a predicted and recorded human trajectory often misrepresent the relative accuracy of trajectories. (3) Since driving involves interactions among multiple agents, evaluation must ideally be interactive, e.g., in simulation. Unfortunately, existing simulators with synthetic sensor data exhibit a significant domain gap to real-world driving. (4) Besides, the lack of a standardized evaluation setup has led to subtle inconsistencies between metrics in existing work, leading to unfair comparisons and inaccurate conclusions [50, 32]. Collectively, these problems hinder progress in the development of AVs, emphasizing the need for more principled benchmarks.
@@ 2 Related Work
Open-Loop Benchmarking with Displacement Errors. Open-loop evaluation protocols commonly measure displacement errors between trajectories of a recorded expert (i.e., of a human driver) and a motion planner. However, several issues concerning evaluation with displacement errors have surfaced recently, particularly on the nuScenes dataset [5]. Given that nuScenes does not provide standardized planning metrics, prior work relied on independent implementations, which led to inconsistencies when reporting or comparing results [50, 32]. Next, most planning models in nuScenes receive the human trajectory endpoint as a discrete direction command [20, 21, 25, 32, 50], thereby leaking ground-truth information into inputs. Moreover, about 75% of the scenarios in nuScenes involve trivial straight driving [32], leading to simple solutions when extrapolating the ego-motion. For instance, AD-MLP demonstrates that an MLP on the kinematic ego status (ignoring perception completely) can achieve state-of-the-art displacement errors [56]. Such blind agents are undeniably dangerous, which highlights a broader concern: displacement metrics are not correlated to closed-loop driving [14, 17, 3, 16]. In this work, we address prevalent issues of nuScenes and propose a standardized driving benchmark with challenging scenarios and an official evaluation server. We derive a navigation goal from the lane graph instead of the human trajectory to prevent label leakage, and propose principled simulation-based metrics as an alternative to displacement errors.
@@ 3 NAVSIM: Non-Reactive Autonomous Vehicle Simulation
Task description. Driving agents in NAVSIM must plan a trajectory, defined as a sequence of future poses, over a horizon of $h$ seconds. Their input contains streams of past frames from onboard sensors, such as cameras, LiDAR, as well as the vehicle’s current speed, acceleration, and navigation goal, jointly termed the ego status. For compatibility with prior work [20, 21, 25, 50], we provide the navigation goal as a one-hot vector with three categories: left, straight, or right.
Non-Reactive Simulation. Traditional closed-loop benchmarks normally infer planners at high frequencies, e.g., 10Hz [17, 26]. However, this requires efficient simulation of all input modalities for the driving agent, including high-dimensional sensor streams in the case of sensor-based approaches. To sidestep this, the core idea of NAVSIM is to evaluate driving agents using a non-reactive simulation. This means driving agents are only queried in the initial frame of each scene. Afterwards, the planned trajectory is kept fixed for the entire trajectory duration. Over this short horizon, no environmental feedback is provided to the driving agent, and the NAVSIM evaluation is purely based on the initial real-world sensor sample. This makes the agent’s task more challenging, limiting simulations to short horizons. We select a horizon of $h=4$ seconds, which has been shown in prior work to be adequate for closed-loop planning [16]. Despite this limitation, non-reactive simulation offers a key advantage: unlike traditional open-loop benchmarks, which mainly compare the planned trajectory to the human driver’s trajectory in a similar setting, it enables the use of simulation outcomes to compute metrics reflecting safety, comfort, and progress. An LQR controller [28] is applied at each simulation iteration to calculate steering and acceleration values, and a kinematic bicycle model [37] propagates the ego vehicle. We execute this pipeline at $10$Hz over the $4$s trajectory horizon. In Sec. 4.1, we show that despite our simplifying assumption, our evaluation results in a much better alignment with closed-loop metrics than traditional open-loop metrics achieve.
@@ 3.1 Generating Standardized and Challenging Train and Test Splits
Dataset. The NAVSIM framework is agnostic to the choice of driving dataset. We choose OpenScene [15], a redistribution of nuPlan [26], the largest annotated public driving dataset. OpenScene includes 120 hours of driving at a reduced frequency of $2$Hz typically considered by end-to-end planning algorithms, resulting in a $90\%$ reduction of data storage requirements compared to nuPlan from over 20 TB to 2 TB. Our agent input, based on OpenScene, comprises eight cameras, each with a resolution of $1920\times 1080$ pixels, and a merged LiDAR point cloud from five sensors. The input includes the current time-step and optionally 3 past frames, totaling $1.5$s at $2$Hz. In principle, any driving dataset that provides annotated HD maps, object bounding boxes, and sensor data can be converted into this format and thus be used with NAVSIM.
@@ 4 Experiments
Figure 3: Closed-Loop Alignment. (a) For each planner, we show open-loop metrics (OLS, PDMS) together with the corresponding closed-loop score (CLS). The trendlines depicting correlations are fit linearly to all (learned and rule-based) planners. Moreover, we analyze different (b) CLS durations $d$, (c) planning frequencies $f$, (d) PDMS horizons $h$, and (e) closed-loop background agent behaviors.
@@ 4.1 Alignment Between Open-Loop and Closed-Loop Evaluation
Due to the heavy computational requirements of closed-loop simulation, we evaluate on the navmini split. This is a new split we create for rapid testing, with 396 scenarios in total that are independent of both navtrain and navtest but filtered using the same strategy (Section 3.1) and hence similarly distributed. We note that nuPlan offers two kinds of background agents: reactive agents along lane centers based on the Intelligent Driver Model (IDM) [48], and non-reactive agents replayed from the dataset, which we employ unless otherwise stated. While reactive simulations of longer or dynamic lengths are generally desirable, e.g. to evaluate long-term decisions, enabling this requires dedicated solutions to long-horizon simulation that are not currently available in nuPlan [13]. Therefore, we default to a fixed closed-loop simulation duration of $d=15\textrm{s}$, and a planning frequency of $f=10\textrm{Hz}$, which are the standard closed-loop simulation settings in nuPlan [26].
@@
Results. The alignment between metrics is presented in Fig. 3 (a-e). Compared to OLS, we consistently observe better closed-loop correlation for PDMS, in terms of Spearman’s (rank) and Pearson’s (linear) correlation coefficients. As shown in (a), PDMS can capture the closed-loop properties of both learned and rule-based planners, whereas distance-based open-loop metrics show a clear misalignment. Decreasing the CLS duration in (b) from $d=15$s to $d=4$s further raises the correlation of PDMS and OLS, as the simulation horizon more closely matches the open-loop counterparts. Interestingly, we observe a higher correlation of open-loop metrics in (c) when reducing the planning frequency to $2$Hz. We expect a lower planning frequency to mitigate cumulative errors and enhance the controller’s stability in simulation, leading to more precise trajectory execution. Moreover, we observe an increase in correlation for longer PDMS horizons in (d), ranging from $h=2$s to $h=8$s. While predicting the future motion over 8s is challenging in uncertain scenarios, our results indicate the value of long horizons when evaluating motion planners. Lastly, replacing the non-reactive background agents with reactive IDM vehicles during closed-loop simulation in (e) has little effect on the correlation, possibly due to the similar difficulty of both tasks [16].
The imbalanced distribution of different types of planners in our study may introduce biases into the overall correlations presented in Fig. 3. To address this, we visualize the individual correlations of each planner type in Fig. 4. The correlation values vary depending on metric range and variance of each planner type. Nevertheless, when examining each type individually, the PDMS is better correlated to the CLS than the OLS, and is always positively correlated.
Figure 4: Planner-Level Alignment of Metrics. We report the correlation coefficients between open-loop metrics (OLS, PDMS) and the closed-loop score (CLS) for the five planner types considered in our study. The PDMS is better correlated to the CLS for every planner type.
@@ 4.2 Analysis of the State of the Art in End-to-End Autonomous Driving
Methods. As a lower bound, we consider the (1) Constant Velocity baseline detailed in Section 3.1. We include an (2) Ego Status MLP as a second "blind" agent, which leverages an MLP for trajectory prediction given only the ego velocity, acceleration and navigation goal. As an established architecture on CARLA, we evaluate our reimplementation of (3) TransFuser [12], which uses three cropped and downscaled forward-facing cameras, concatenated into a $1024\times 256$ image, and a rasterized BEV LiDAR input for predicting waypoints. It performs 3D object detection and BEV semantic segmentation as auxiliary tasks. We then consider (4) Latent TransFuser (LTF) [12], which shares the same architecture as TransFuser but replaces the LiDAR input with a learned embedding, hence requiring only camera inputs. Moreover, we provide two state-of-the-art end-to-end architectures for open-loop trajectory prediction on nuScenes. (5) UniAD [21] incorporates a wide range of tasks, such as mapping, tracking, motion, and occupancy prediction in a semi-sequential architecture, which processes feature representations through several transformer decoders culminating in a trajectory planning module. (6) PARA-Drive [50] uses the same auxiliary tasks, but parallelizes the network architecture, such that the auxiliary task heads are trained in parallel with a shared encoder. Both UniAD and PARA-Drive use a BEVFormer backbone [31], which encodes the eight surround-view $1920\times 1080$ camera images over four temporal frames into a BEV feature representation. Implementation details for all methods are provided in the supplementary material.
@@
Method PDMS $\uparrow$
TransFuser [12] $83.9\pm 0.4$
LTF [12] $83.5\pm 0.6$
Ego Status MLP $66.4\pm 0.9$
Hydra-MDP [30] $91.3$
Constant Velocity $20.6$
Table 3: Leaderboard 1.1.
NAVSIM 1.1 Leaderboard. Due to the lasting interest after the challenge, we re-opened a public evaluation server using navtest as the evaluation split. The leaderboard encourages multi-seed submissions and includes reproducibility requirements for openly releasing code and model weights. We populated the leaderboard with 3 training seeds of our learned baselines, as shown in Table 3. For reference, we also include a single seed of the 2024 challenge winner [30] and constant velocity baseline. Further information is provided in the supplementary material and leaderboard webpage [^2] https://huggingface.co/spaces/AGC2024-P/e2e-driving-navsim.
@@ 5 Discussion
We present NAVSIM, a framework for non-reactive AV simulation. We address shortcomings of existing driving benchmarks and propose standardized but configurable simulation-based metrics for benchmarking driving policies. For accessibility, we provide challenging scenario splits and simple data curation methods. We show that our evaluation protocol is better aligned to closed-loop driving, benchmark an established set of end-to-end planning baselines from CARLA and nuScenes, and present the results of our inaugural competition. We hope that NAVSIM can serve as an accessible toolkit for AV researchers that bridges the gap between simulated and real-world driving.
Need for Reactive Simulation. While we show improvements over displacement errors, several aspects of driving remain unaddressed by evaluation in NAVSIM. A high PDMS does not always imply a high CLS, since our framework does not consider reactiveness or the compounding accumulation of errors in closed-loop simulation. Moreover, as in CLS, rear-end collisions into the ego vehicle are currently not classified as "at-fault", resulting in little importance given to the scene behind the vehicle in NAVSIM. In the future, data-driven sensor or traffic simulation could alleviate these issues, once such methods mature and become computationally tractable. Given these limitations of the current framework, we strongly encourage the use of graphics-based closed-loop simulators, such as CARLA [17], as complementary benchmarks to NAVSIM when developing planning algorithms.
Simplicity of Metrics. As a starting point, NAVSIM offers both interpretable open-loop subscores and a scalarizing function, which lets us provide a final score and ranking for participants in the challenge. In the future, multi-objective evaluation and other aggregation functions might be required. Moreover, closed-loop metrics also face problems, i.e., PDMS inherits several weaknesses of nuPlan’s CLS. Both scores do not regard certain traffic rules (e.g., stop-sign or traffic light compliance) or concepts such as transit and fuel efficiency. In the future, we aim to improve the subscore definitions (e.g. the at-fault collision logic) and add more subscores during aggregation.
Call for Datasets. Certain limitations of the nuPlan dataset persist in NAVSIM, such as missing classes in the label space, minor errors in camera parameters, or noise in vehicle poses and 3D annotations. Our analysis might favor methods that are robust to such inconsistencies. In addition, the lack of road elevation data in our representation presents a challenge for integrating scenarios based on 3D map annotations. We aim to support more datasets in the future, and advocate for more open dataset releases by the community for accelerating progress in autonomous driving.